. Thirty-six percent of all U. S. households have a pet dog. Thirty-four percent of all U. S. households have a pet cat. Two percent of all U. S. households have a horse and interestingly, eighty percent of households with a horse, also own a dog. (Of course, households with a horse and a dog would also be part of the 36% with a dog.) A U. S. household is chosen at random.

(a)  What is the probability it has a dog?

 

(b)  What is the probability it has a dog or a cat?

 

(c)  What is the probability it has a horse?

 

(d)  What is the probability it has a dog and a horse?

 

(e)  What is the probability it has a horse only?

 

(f)   What is the probability it has a dog only?

 


1
Expert's answer
2021-08-10T07:27:16-0400

Let use the sample of 1000 U.S. households.

Number of households that have a pet dog "= 1000 \times 0.36 = 360"

Number of households that have a pet cat "= 1000 \times 0.34 = 340"

Number of households that have a horse "= 1000 \times 0.02=20"

Number of households with a horse, also own a dog "= 20 \times 0.8 = 16"

(a) P(has a dog) "= \frac{360}{1000}= 0.36"

(b) P(has a dog or a cat) "= \frac{360+340}{1000}= 0.7"

(c) P(has a horse) "= \frac{20}{1000}= 0.02"

(d) P(has a dog and a horse) "= \frac{16}{1000}=0.016"

(e) P(has a horse only) "= \frac{20-16}{1000} = 0.004"

(f) P(has a dog only) "= \frac{360-16}{1000}=0.344"